a.    What is the probability of the first selected person is a left-handed men and the second selected person is a right-handed men?

 


1
Expert's answer
2020-08-02T17:56:15-0400

suppose there are 10 persons and 4 are left handed and 6 right handed persons.


Case 1: without replacement


P(first is left handed and second right handed)= "\frac{4}{10}*\frac{6}{9}" = "\frac{24}{90}" = 0.27


Case 2: With replacement

P(first is left handed and second right handed)= "\frac{4}{10}*\frac{6}{10}" ="\frac{24}{100}" =0.24